OpenStudy (anonymous):

Solve using the formula for the future value of an ordinary annuity. Find the value of the annuity at the end of the indicated time period. The amount, frequency of deposits, annual interest rate, and time period are given. Do not round until the final answer. Then, round to the nearest cent. Amount: $225; quarterly; 9%; 8 years.

terenzreignz (terenzreignz):

Oh my :> Were you actually given the said formula? ^_^

terenzreignz (terenzreignz):

Hey @brimichelle15 don't die on me :3

OpenStudy (anonymous):

\[\Large A=\frac{P \left(\left(\frac{r}{n}+1\right)^{ n t}-1\right)}{\frac{r}{n}} \]

OpenStudy (anonymous):

In your case P=225 r=.09 t=8

OpenStudy (anonymous):

n=4